OpenStudy (niahloves):

The perimeter of this shape is 77 cm. What is the length of side d?

OpenStudy (anonymous):

Could you make a drawing?

OpenStudy (jameshorton):

i did this one

OpenStudy (anonymous):

21.4 = D

OpenStudy (niahloves):

The length of a rectangle is 5 times the width. What is the perimeter of the rectangle if the width is 12 inches? A. 72 inches B. 120 inches C. 144 inches D. 240 inches

OpenStudy (niahloves):

@Firez

OpenStudy (niahloves):

@superman38

OpenStudy (anonymous):

c

OpenStudy (niahloves):

How many feet are there in 456 in.

OpenStudy (jameshorton):

38

OpenStudy (anonymous):

38

OpenStudy (niahloves):

In which measurement situations would the most appropriate unit of measurement be meters? Choose exactly two answers that are correct. A. width of a finger B. height of a tree C. depth of a well D. distance from the earth to the sun

OpenStudy (anonymous):

b nd c

OpenStudy (niahloves):

one more

OpenStudy (niahloves):

(1 pt) A rectangle has a length of 10 cm and a perimeter of 46 cm. Which equation could be used to find the width of the rectangle? A. 10w = 46 B. w + 10 = 46 C. 2w + 20 = 46 D. 2w + 10 = 46

OpenStudy (anonymous):

b nd c

OpenStudy (niahloves):

i have to pick one

OpenStudy (jameshorton):

c

OpenStudy (anonymous):

c